OpenStudy (anonymous):

Let f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (austinl):

For this you just plug in -1 for all instances of x in f(x).

OpenStudy (anonymous):

just plug -1 into your x values so F(x)= (-1)2 - 8(-1) +5 so F(x)= -2 -(-8) +5 when subtracting a negative number you add it so the question becomes F(x)= -2 +8 +5 f(-1)=11

OpenStudy (austinl):

\(f(x)=x^2-8x+5\\ f(-1)=(-1)^2-8(-1)+5\)

OpenStudy (anonymous):

unless that's x^2?

OpenStudy (anonymous):

if it is the equation turns to F(-1)= 1 + 8 +5

OpenStudy (anonymous):

it x^2